# 0.5 eV QCD Axion Dark Matter

### v44

@@ 3.2 Big Bang Temperature from QCD Axion Angle Jump $\Delta\theta_{A}=\pi$
In conclusion, we can simply replace the $\approx$ symbol in Eq. (3.6) with an $=$ sign. The cosmological derivation of the rest-mass energy of the 0.5 eV QCD axion dark matter is now complete. The key step was to use the fact that there are twelve flavors of 0.5 eV QCD axion dark matter to deduce that there are six such axions for every photon in the CMB.
@@ B $432\pm 1$ [63]
$I$ $4079.57\pm 0.18$ $\nu_{A\theta}(6)-\nu_{A\theta}(4)$
$II$ $4084.6\pm 1.0$ $\nu_{A\theta}(2)-\nu_{\theta}(A)$
$III$ $4079.6\pm 3.0$ $\nu_{A\theta}(3)-\nu_{\theta}(C)$
$IV$ $4075.1\pm 3.2$ $\nu_{A\theta}(5)-\nu_{\theta}(E)$
Table 3: Axion Compton Frequency Determinations— The Compton frequency of the axion $\nu_{A}=m_{A}c/h$ is determined by four independent analyses. The independent but consistent determinations are based on the polystyrene axion resonance frequencies $\nu_{A\theta}$ from Table 1 and the axion-active polystyrene phonon frequencies $\nu_{\theta}$ from Table 2. Determination $I$ gives the rest-mass of the axion $m_{A}(I)=h\nu_{A}(I)/c=(0.50580\pm 0.00002)~{}{\rm eV/c^{2}}$ to 44 parts per million precision.
@@
In the case of photons, the coupling to 0.5 eV QCD axion dark matter from the local dark matter halo shows up as a change in the effective value of the electromagnetic “fine-structure” constant $\alpha$ that governs quantum electrodynamic (QED) effects on charged leptons [67]. To leading order in $\alpha\approx 1/137$, QED effects — such as the shift $a_{\ell}=(g_{\ell}-2)/2\approx\alpha/(2\pi)=1.164\times 10^{-3}$ of the magnetic dipole moment of the charged lepton $\ell$ — come from an emergent mass term in the Dirac equation created by the proper electromagnetic radiation field of the charged lepton itself [68]. The axion-photon coupling in Eq. (3.7) changes the photon propagator that enters this emergent mass term.
Effectively, the change in the photon propagator can be absorbed into a change in the value of the fine-structure constant
$\frac{\Delta\alpha}{\alpha}$ $=$ $g_{A\gamma\gamma}m_{A}c^{2}\sqrt{\frac{\rho_{A}N_{\ell}V_{\ell}T_{\ell}}{\hbar}},$ (5.2)
@@
where $N_{\ell}$ is the number of charged leptons $\ell$ interacting with the proper radiation field, $V_{\ell}$ is the luminous volume where this interaction takes place, and $T_{\ell}$ is the coherent life-time of the charged lepton within the luminous volume. Using $\rho_{A}=\rho_{\rm DM}=(0.30\pm 0.03)~{}{\rm GeV/cm^{3}}$ for the local energy density of 0.5 eV QCD axion dark matter, we can predict the size of the resulting shift in the muon magnetic dipole moment measured with muon spin precession in magnetic storage rings
$\Delta a_{\mu}$ $=$ $a_{\mu}\frac{\Delta{\alpha}}{\alpha}=(170\pm 10)\times 10^{-11}\left(\frac{N_{\mu}}{2\times 10^{3}}\right)^{1/2}\left(\frac{V_{\mu}}{3\times 10^{4}~{}{\rm cm^{3}}}\right)^{1/2}\left(\frac{T_{\mu}}{64~{}{\rm\mu sec}}\right)^{1/2}$ (5.3)
@@
where $N_{\mu}$ is the number of muons whose spins are precessing within the magnetic storage ring, $V_{\mu}=2\pi^{2}\sigma_{x}\sigma_{y}R$ is the volume of the muon beam of width $\sigma_{x}$ and height $\sigma_{y}$ within the storage ring of radius $R$, and $T_{\mu}=\tau_{\mu}\sqrt{(p_{\mu}/(m_{\mu}c)^{2}+1}$ is the life-time of the muon beam with momentum $p_{\mu}$ in the rest-frame of the ring for muon rest-mass $m_{\mu}=105.7~{}{\rm MeV/c^{2}}$ [23] and rest-frame lifetime $\tau_{\mu}=2.20~{}{\rm\mu sec}$ [23]. The predicted shifts are shown in Table 4 together with the input parameters for the three muon $g-2$ experiments whose beam parameters are given in Table 5.
@@
Table 4: 0.5 eV QCD Axion Dark Matter Shifts Muon $g-2$ The number of muons $N_{\mu}$ stored within luminous volume $V_{\mu}$ with life-time $T_{\mu}$ for Brookhaven (BNL E821) [69], Fermilab (FNAL E989) [70], and J-PARC (JPARC E34) [71] muon spin precession experiments in magnetic storage rings with beam parameters in Table 5 yield a percent-level prediction for the shift $\Delta a_{\mu}=\Delta g_{\mu}/2\approx\Delta\alpha/(2\pi)$ of the muon magnetic dipole moment $a_{\mu}=(g_{\mu}-2)/2$ by 0.5 eV QCD axion dark matter from the local dark matter “halo” of the galaxy given by Eq. (5.3). J-PARC muon number is the design goal. Brookhaven and Fermilab muon numbers are measured values [72].
@@
ATLAS LHC Run 1 $pp$ 6.24 [76] $13.1\pm 0.6$
CMS LHC Run 2 $pp$ 6.24 [76] $6.7\pm 0.3$
JADE,TASSO PETRA $e\overline{e}$ 35 [77] $-390\pm 17$
Table 6: 0.5 eV QCD Axion Dark Matter Shifts $\alpha_{s}$ Experiments at high-energy colliders of type proton-antiproton ($p\overline{p}$), proton-proton ($pp$) and electron-positron ($e\overline{e}$) with vector boson momentum $Q$ in the lab frame. The predicted shift of the strong coupling constant due to the coupling of gluons to the 0.5 eV QCD axion dark matter from the local dark matter halo given by Eq. (5.6) is $\Delta\alpha_{s}(M_{W})$ for $W$ boson production from quark-antiquark annihilation during hadron collisions ($pp$ and $p\overline{p}$). The shift is $\Delta\alpha_{s}(M_{Z})$ for quark-antiquark dijet production from electron-positron annihilation ($e\overline{e}$).
@@
Standard Model [83] $80,356~{}(5)$ 0 $80,356~{}(5)(0)$
ATLAS [84, 85] $80,367~{}(16)$ $13.3~{}(6)$ $80,354~{}(16)(0.6)$
CDF [81] $80,433.5~{}(9.4)$ $78~{}(4)$ $80,355.5~{}(9.4)(4)$
CMS [86] $80,360.2~{}(9.9)$ $6.8~{}(3)$ $80,353.4~{}(9.9)(0.3)$
@@
Table 9: 0.5 eV QCD Axion Dark Matter Shifts $a_{\mu}^{\pi\pi}$eading order hadronic contribution $a_{\mu}^{\pi\pi}$ to the muon magnetic dipole moment determined by pion pair production $\pi^{+}\pi^{-}(\gamma)$ from electron-positron annihilation with center-of-mass energy in the range $0.6\leq\sqrt{s}\leq 0.88~{}{\rm GeV}$ at various experiments compiled by ref. [87]. The luminous four-volume created by the colliding electron and positron bunches produces the shift $\Delta a_{\mu}^{\pi\pi}$ given by Eq. (5.15). The Standard Model value $a_{\mu}^{\pi\pi}({\rm SM})=a_{\mu}^{\pi\pi}+\Delta a_{\mu}$ corrects for the effect of 0.5 eV QCD axion dark matter and agrees between different experiments yielding a determination to 0.4% (See Eq. (5.16)).
@@ 6 Conclusion
The main effect of 0.5 eV QCD axion dark matter on Big Bang nucleosynthesis is shown in Fig. 2 ([17, 97], ref. [51], pg. 184–196). In the standard expanding universe cosmology, there are two key moments in the early universe’s expansion history that are critical to the efficiency of converting nucleons — neutrons $n$ and protons $p$— into the nuclei of the light chemical elements — primarily alpha particles $\alpha$ with a tiny bit of deuterons $d$. The domination of the early expansion history of the universe by 0.5 eV QCD axion dark matter shrinks the gap in time between these two key moments compared to the standard expanding universe cosmology resulting in a significant increase in $\alpha$ production for a given amount of nucleons.

### v45

@@ 3.2 Big Bang Temperature from QCD Axion Angle Jump $\Delta\theta_{A}=\pi$
In conclusion, we can simply replace the $\approx$ symbol in Eq. (3.6) with an == sign. The cosmological derivation of the rest-mass energy of the 0.5 eV QCD axion dark matter is now complete. The key step was to use the fact that there are twelve flavors of 0.5 eV QCD axion dark matter to deduce that there are six such axions for every photon in the CMB.
@@ B $432\pm 1$ [63]
II $4079.57\pm 0.18$ $\nu_{A\theta}(6)-\nu_{A\theta}(4)$
IIII $4084.6\pm 1.0$ $\nu_{A\theta}(2)-\nu_{\theta}(A)$
IIIIII $4079.6\pm 3.0$ $\nu_{A\theta}(3)-\nu_{\theta}(C)$
IVIV $4075.1\pm 3.2$ $\nu_{A\theta}(5)-\nu_{\theta}(E)$
Table 3: Axion Compton Frequency Determinations— The Compton frequency of the axion $\nu_{A}=m_{A}c/h$ is determined by four independent analyses. The independent but consistent determinations are based on the polystyrene axion resonance frequencies $\nu_{A\theta}$ from Table 1 and the axion-active polystyrene phonon frequencies $\nu_{\theta}$ from Table 2. Determination II gives the rest-mass of the axion $m_{A}(I)=h\nu_{A}(I)/c=(0.50580\pm 0.00002)~{\rm eV/c^{2}}$ to 44 parts per million precision.
@@
In the case of photons, the coupling to 0.5 eV QCD axion dark matter from the local dark matter halo shows up as a change in the effective value of the electromagnetic “fine-structure” constant $\alpha$ that governs quantum electrodynamic (QED) effects on charged leptons [67]. To leading order in $\alpha\approx 1/137$, QED effects — such as the shift $a_{\ell}=(g_{\ell}-2)/2\approx\alpha/(2\pi)=1.164\times 10^{-3}$ of the magnetic dipole moment of the charged lepton $\ell$ — come from an emergent mass term, $\int(dx^{\prime})M(x,x^{\prime})\psi(x^{\prime})$, in the Dirac equation for the lepton wavefunction, $\psi(x)$, with mass operator, $M(x,x^{\prime})$, given, to second order in $e$, by the following expression [68]:
$M(x,x^{\prime})=m_{0}\delta(x-x^{\prime})+ie^{2}\gamma_{\mu}G(x,x^{\prime})\gamma_{\nu}D_{+}(x-x^{\prime}),$ (5.2)
where $G(x,x^{\prime})$ is the charged lepton Green’s function in the presence of the external electromagnetic field, and $m_{0}$ is the “bare” rest-mass energy of the charged lepton. The second term is created by the proper electromagnetic radiation field of the charged lepton itself, and the axion-photon coupling in Eq. (3.7) changes the photon propagator, $D_{+}(x-x^{\prime})$, that enters this emergent mass term to create the axion-induced change, $\Delta M(x,x^{\prime})$, in the mass operator, as follows:
$\Delta M(x,x^{\prime})=ie^{2}g_{A\gamma\gamma}\tilde{\phi}_{A}(p_{A})\epsilon_{\mu\nu\lambda\sigma}(-ip_{A}^{\mu})\gamma^{\nu}G(x,x^{\prime})\gamma^{\sigma}\partial^{\lambda}D_{+}(x-x^{\prime})$ (5.3)
where, $\tilde{\phi}_{A}(p_{A})=\sqrt{\rho_{A}NVT}/(m_{A}c^{2})$ is the Fourier amplitude of the axion dark matter from the local “halo” of the galaxy with energy density $\rho_{A}$ within the effective luminous four-volume $NVT$ created by the $N$ charged leptons in spatial volume $V$ during the coherence time $T$ where the interaction with the proper radiation field takes place.
Following a straightforward analysis along the lines of the derivation of the leading order result $a_{\ell}=\alpha/2\pi$ from the mass operator $M(x,x^{\prime})$ in Eq. (5.2) [68], one finds the change in the mass operator $\Delta M(x,x^{\prime})$ in Eq. (5.2) leads to the change, $\Delta a_{\ell}=\Delta\alpha/2\pi$, with the effective shift in the fine-structure constant, $\Delta\alpha$, given by
$\Delta\alpha$ $=$ $\alpha g_{A\gamma\gamma}m_{A}c^{2}\sqrt{\frac{\rho_{A}N_{\ell}V_{\ell}T_{\ell}}{\hbar}},$ (5.4)
@@
Using $\rho_{A}=\rho_{\rm DM}=(0.30\pm 0.03)~{\rm GeV/cm^{3}}$ for the local energy density of 0.5 eV QCD axion dark matter, we can predict the size of the resulting shift in the muon magnetic dipole moment measured with muon spin precession in magnetic storage rings
$\Delta a_{\mu}$ $=$ $\frac{\Delta{\alpha}}{2\pi}=(170\pm 10)\times 10^{-11}\left(\frac{N_{\mu}}{2\times 10^{3}}\right)^{1/2}\left(\frac{V_{\mu}}{3\times 10^{4}~{\rm cm^{3}}}\right)^{1/2}\left(\frac{T_{\mu}}{64~{\rm\mu sec}}\right)^{1/2}$ (5.5)
@@
where $N_{\mu}$ is the number of muons whose spins are precessing within the magnetic storage ring, $V_{\mu}=2\pi^{2}\sigma_{x}\sigma_{y}R$ is the volume of the muon beam of width $\sigma_{x}$ and height $\sigma_{y}$ within the storage ring of radius $R$, and $T_{\mu}=\tau_{\mu}\sqrt{(p_{\mu}/(m_{\mu}c)^{2}+1}$ is the life-time of the muon beam with momentum $p_{\mu}$ in the rest-frame of the ring for muon rest-mass $m_{\mu}=105.7~{\rm MeV/c^{2}}$ [23] and rest-frame lifetime $\tau_{\mu}=2.20~{\rm\mu sec}$ [23]. The predicted shifts are shown in Table 4 together with the input parameters for the three muon g2g-2 experiments whose beam parameters are given in Table 5.
@@
Table 4: 0.5 eV QCD Axion Dark Matter Shifts Muon g2g-2 The number of muons $N_{\mu}$ stored within luminous volume $V_{\mu}$ with life-time $T_{\mu}$ for Brookhaven (BNL E821) [69], Fermilab (FNAL E989) [70], and J-PARC (JPARC E34) [71] muon spin precession experiments in magnetic storage rings with beam parameters in Table 5 yield a percent-level prediction for the shift $\Delta a_{\mu}=\Delta g_{\mu}/2\approx\Delta\alpha/(2\pi)$ of the muon magnetic dipole moment $a_{\mu}=(g_{\mu}-2)/2$ by 0.5 eV QCD axion dark matter from the local dark matter “halo” of the galaxy given by Eq. (5.5). J-PARC muon number is the design goal. Brookhaven and Fermilab muon numbers are measured values [72].
@@
ATLAS LHC Run 1 pppp 6.24 [76] $13.1\pm 0.6$
CMS LHC Run 2 pppp 6.24 [76] $6.7\pm 0.3$
JADE,TASSO PETRA $e\overline{e}$ 35 [77] $390\pm 17$
Table 6: 0.5 eV QCD Axion Dark Matter Shifts $\alpha_{s}$ Experiments at high-energy colliders of type proton-antiproton ($p\overline{p}$), proton-proton (pppp) and electron-positron ($e\overline{e}$) with vector boson momentum $Q$ in the lab frame. The predicted shift of the strong coupling constant due to the coupling of gluons to the 0.5 eV QCD axion dark matter from the local dark matter halo given by Eq. (5.8) is $\Delta\alpha_{s}(M_{W})$ for $W$ boson production from quark-antiquark annihilation during hadron collisions (pppp and $p\overline{p}$). The shift is $\Delta\alpha_{s}(M_{Z})$ for quark-antiquark dijet production from electron-positron annihilation ($e\overline{e}$).
@@
Standard Model [83] 80,356(5)80,356~(5) 0 80,356(5)(0)80,356~(5)(0)
ATLAS [84, 85] 80,367(16)80,367~(16) 13.3(6)13.3~(6) 80,354(16)(0.6)80,354~(16)(0.6)
CDF [81] 80,433.5(9.4)80,433.5~(9.4) 78(4)78~(4) 80,355.5(9.4)(4)80,355.5~(9.4)(4)
CMS [86] 80,360.2(9.9)80,360.2~(9.9) 6.8(3)6.8~(3) 80,353.4(9.9)(0.3)80,353.4~(9.9)(0.3)
@@
Table 9: 0.5 eV QCD Axion Dark Matter Shifts $a_{\mu}^{\pi\pi}$Leading order hadronic contribution $a_{\mu}^{\pi\pi}$ to the muon magnetic dipole moment determined by pion pair production $\pi^{+}\pi^{-}(\gamma)$ from electron-positron annihilation with center-of-mass energy in the range $0.6\leq\sqrt{s}\leq 0.88~{\rm GeV}$ at various experiments compiled by ref. [87]. The luminous four-volume created by the colliding electron and positron bunches produces the shift $\Delta a_{\mu}^{\pi\pi}$ given by Eq. (5.15). The Standard Model value $a_{\mu}^{\pi\pi}({\rm SM})=a_{\mu}^{\pi\pi}+\Delta a_{\mu}$ corrects for the effect of 0.5 eV QCD axion dark matter and agrees between different experiments yielding a determination to 0.4% (See Eq. (5.18)).
@@ 6 Conclusion
The main effect of 0.5 eV QCD axion dark matter on Big Bang nucleosynthesis is shown in Fig. 2 ([17, 97], ref. [51], pg. 184–196). In the standard expanding universe cosmology, there are two key moments in the early universe’s expansion history that are critical to the efficiency of converting nucleons — neutrons $n$ and protons pp— into the nuclei of the light chemical elements — primarily alpha particles $\alpha$ with a tiny bit of deuterons $d$. The domination of the early expansion history of the universe by 0.5 eV QCD axion dark matter shrinks the gap in time between these two key moments compared to the standard expanding universe cosmology resulting in a significant increase in $\alpha$ production for a given amount of nucleons.